A firm operates two plants. Plant 1 as a MC = 4q1, and plant 2 as a MC = 2q2. Total demand is 150 units. Firm is currently producing 40 units out of plant 1 and 110 units out of plant 2. Is the firm maximizing profits? Why or why not? What would you do differently (if anything)?

What is missing?
Answer #1

At eqm, MC of Production is Equalized across all plants,

Now MC1= 4*40 = 160

MC2 = 2*110 = 220

So MC is not Equalized .

So firm is not Maximizing profits .

So at π Maximization,

MC1 = MC2

4q1 = 2q2

2q1 = q2 & q1+q2 = 150

So , q2 = 150-q1

2q1 = 150-q1

3q1 = 150

q1* = 50, q2* = 150-50 = 100

SO AT EQM,

q1* = 50 & q2* = 100